OpenStudy (anonymous):

The density of cylindrical rod was measured by ρ =4m/ ПD2l The percentage error in m, D, l are 1%, 1.% and 0.5%. Calculate % error in the calculated value of density.

OpenStudy (maheshmeghwal9):

\[\frac{\Delta \rho }{\rho}=(+1)\frac{\Delta m}{m}+ (-1)\frac{\Delta D}{D} + (-1)\frac{\Delta I}{I}\]\[\implies \frac{\Delta \rho}{\rho}= (+1 \times 1)+(-1 \times 1)+(-1 \times 0.5) \space percent\] => ( 1-1-0.5 ) % => (-0.5) % Answer.

OpenStudy (maheshmeghwal9):

Am I right?

OpenStudy (maheshmeghwal9):

I mean wt is the answer?

OpenStudy (anonymous):

That's the problem, i don't have an answer.. but know what ? i got 4.5% as answer!

OpenStudy (maheshmeghwal9):

how u gt?

OpenStudy (anonymous):

Note: The above method can be applied only if the percentages are less than 5. That's cause it's an approximation.

OpenStudy (maheshmeghwal9):

thanx for telling i didn't know that upto today.

OpenStudy (maheshmeghwal9):

so wt for more than 5?

OpenStudy (anonymous):

Do it the long way. That is, write the expression for density got using values with error. write expn for actual density. relate all the respective quantities using the error %ges.

OpenStudy (maheshmeghwal9):

ok!